# Can industrial overcapacity enable seasonal flexibility in electricity use? A case study of aluminum smelting in China

### 26 Mar 2026 v2

@@ Abstract
In many countries, declining demand in energy-intensive industries (EIIs) such as cement, steel, and aluminum is leading to industrial overcapacity. Although industrial overcapacity is traditionally envisioned as problematic and resource-wasteful, it could unlock EIIs’ flexibility in electricity use. Here, using China’s aluminum smelting industry as a case study, we evaluate the system-level cost-benefit of retaining EII overcapacity for flexible electricity use in decarbonized energy systems. We find that overcapacity can enable aluminum smelters to adopt a seasonal operation paradigm, ceasing production during winter load peaks that are exacerbated by heating electrification and renewable seasonality. This seasonal operation paradigm could reduce the investment and operational costs of China’s decarbonized electricity system by 23-32 billion CNY/year (11-15% of the aluminum smelting industry’s product value), sufficient to offset the increased smelter maintenance and product storage costs associated with overcapacity. It may also provide an opportunity for seasonally complementary labor deployment across the aluminum smelting and thermal power generation sectors, offering a potential pathway for mitigating socio-economic disruptions caused by industrial restructuring and energy decarbonization.
@@ 1 Introduction
In this paper, we aim to understand the value of aluminum smelting overcapacity for providing demand flexibility in China’s decarbonized energy system. We examine these effects across both physical infrastructure and socio-economic dimensions (e.g., labor deployment), as both are essential for charting a sustainable path for the aluminum industry. However, the computational complexity of modeling and optimizing the intricate coupling between two already large and complex systems—the electricity system and aluminum smelters— is a significant barrier to such analysis and has not been done before. We developed an integrated provincial-level power-smelting co-optimization model (Figure 1). This approach couples the PyPSA power system framework [6] with a detailed aluminum smelter operation model [47] for hourly co-optimization. To solve this computationally heavy 8760-hour simulation, we adopted a Data-Driven Dimension Reduction (D3R) framework [37] paired with an iterative decomposition algorithm. First, D3R linearizes complex smelter constraints and generates an initial co-optimized solution to make the power system model easier to solve. The power system model then passes hourly locational marginal prices (LMPs) to the smelters. Using these price signals, each smelter independently optimizes its discrete production decisions (e.g., shutdown and restart costs) in parallel. This price-load feedback loop iterates until convergence, bridging modeling high-fidelity smelting physics with large-scale grid planning (see Section 4 for details).
Figure 1: Schematic diagram of the integrated power-aluminum co-optimization framework. To overcome the computational intractability of an 8760-hour horizon, the PyPSA-based grid model and detailed smelter operations are decoupled and solved iteratively through the exchange of price signals (LMPs) and industrial load profiles, aided by the D3R framework.
@@
2 Results
@@
The restart-up costs are measured from the electricity input side. Exchange rates throughout the study (annual averages 2020): 1 USD = 6.900 CNY; 1 EUR = 7.868 CNY; 1 USD = 0.877 EUR. The electricity system cost data is primarily from the PyPSA-China model [65], with the specific Chinese cost assumptions and their original sources documented in Supplementary Note 4.4. Long-duration storage includes H2 and pumped hydro storage. Our core scenario is defined by the combination of Mid primary aluminum demand, Mid smelter operational flexibility, Mid technology costs, and Inflexible labor.
This study employs a scenario analysis approach to explore the impact of key uncertainties on the system-level value of aluminum smelting overcapacity. As detailed in Table 1, we evaluate four primary dimensions: (i) primary aluminum demand, which determines the industry’s aggregate flexibility potential by defining the size of the sector; (ii) smelter operational flexibility, representing the physical minimum load constraints and the economic penalty (e.g., restart costs) of non-baseload operation; (iii) technology costs, reflecting the relative competitiveness of demand-side flexibility against supply-side flexibility alternatives (e.g., batteries and hydrogen); and (iv) labor flexibility, which dictates whether workforce costs can be mitigated during seasonal shutdowns. (see Method 4.1 for more details) To ensure the robustness and traceability of these scenario settings, the bounds for each dimension are systematically derived from empirical literature and industry standards. For technology costs, the “Mid” scenario adopts peer-reviewed trajectories from the PyPSA-China[65] dataset, while the “Low” and “High” bounds for the costs are scaled by 20-20% and +50+50%, respectively, adhering to the AACE International Class 4 estimate guidelines [4].
@@ 2.1 Inevitable overcapacity in aluminum smelting in the recycle era
2.2 Electricity system cost reduction compared to no-overcapacity case
@@
2.3 Cost-benefit analysis of retaining smelting overcapacity
@@
Figure 4: Trade-off analysis and cost-effective strategy for retaining overcapacity. In (a), electricity system cost savings, smelter cost increase, and net system benefit with different retained capacities in 2050 in the core scenario are shown. The maximal net benefit (electricity system cost reduction minus smelter increased costs) is achieved with 30% overcapacity in 2050. In (b), the mean optimal overcapacity rate increases from 6% in 2030 to 30% in 2050, due to higher flexibility requirements as the electricity system approaches net-zero emissions. However, it should be noted that the cost-effective overcapacity of aluminum smelting still decreases across the years due to declining primary aluminum demand. The net benefit across different scenarios first decreases as primary aluminum demand declines to 2040 levels, then increases in the net-zero emissions scenario in 2050, where load flexibility is more valuable.
@@
2.4 Seasonal operation paradigm enabled by overcapacity
@@
(b) Monthly electricity demand and generation in 2050. (Net demand = electricity demand - electricity generation)
Figure 5: Seasonal operation of aluminum smelters is complementary to energy system seasonality. In (a), smelter production largely ceases from mid-November to mid-March with demand met by stored inventory, but operates at full capacity from late March to early November. Comparing (a) with (b), the electricity consumption of aluminum smelting is largely complementary to China’s monthly net electricity demand, acting as a crucial seasonal flexibility resource to mitigate the winter electricity demand peaks driven by heating electrification.
@@
Panel (b) reveals the underlying system driver for this behavior: the emergence of a “seasonal duck curve”. Driven by high renewable penetration and heating electrification, China’s net electricity demand (residual demand after subtracting renewable generation) exhibits a distinct and prolonged winter peak. During these winter periods, renewable energy output, especially from hydropower, is at its minimum while heating demand peaks. Although wind power also increases in the winter, it is outpaced by the surge in heating demand, forcing the power system to rely more heavily on gas and coal generation to ensure power supply reliability. Here, the national aggregate solar output obscures regional differences (e.g., reduced summer solar generation in Yunnan due to the monsoon season). For provincial results, see Supplementary Figure 8.
Figure 6: Comparison of the change in levelized cost per tonne of aluminum between the $20$ baseline and various 2050 overcapacity levels. The composition of the levelized cost per tonne of aluminum is shown across three overcapacity levels in the core scenario and compared against a 2020 baseline. We find that maintaining 30% overcapacity is beneficial for the smelters, reducing aluminum production costs by over 1,500 CNY/tonne (approximately 9% of current production costs), with the drop in electricity bills offsetting the maintenance costs of overcapacity and increased restart, storage, and capital costs.
@@ 3 Discussion
Realizing this seasonal paradigm also requires addressing its socio-economic implications, particularly concerning labor. However, because the labor intensity of modern primary aluminum production is exceptionally low (accounting for only 2%–3% of total costs [49]), the immense energy cost savings achieved during high-generation periods are more than sufficient to cover standby wages or training stipends during winter lulls. As financial incentives make the transition to seasonal operations economically compelling, exploring potential employment complementarity under this paradigm represents a promising avenue for future research and policy design. For instance, the winter production lull in aluminum smelting naturally coincides with the operational peaks of thermal power plants in a decarbonized grid. While facilitating such inter-industry job mobility requires specific spatial planning (e.g., large integrated industrial parks) and targeted institutional support, it is not without precedent; seasonal labor deployment has long been successfully managed in sectors like agriculture and tourism. Ultimately, supported by carefully designed market and policy mechanisms, the seasonal operation of energy-intensive industries holds the potential to unlock broader societal benefits, transforming structural overcapacity into a strategic asset for employment resilience and social stability.
4 Methods
4.1 Data Source and Scenario Settings
@@
In our model, China’s domestic aluminum smelters are required to meet the annual primary aluminum demand, defined as the total demand minus the availability of recycled aluminum. This primary demand is a critical factor determining the system-level value of flexibility and the optimal level of strategically retained overcapacity: intuitively, a higher primary demand implies a larger industrial base with greater cumulative potential for seasonal load shifting and peak shaving. We focus exclusively on primary aluminum smelters for seasonal flexibility, modeling recycled (secondary) aluminum production as a continuous baseload process. This constraint is driven by three techno-economic factors: (1) unlike the legacy overcapacity in primary smelting, the rapidly expanding recycling sector requires substantial new capital investment, necessitating high utilization rates to amortize marginal capital expenditures; (2) the low energy intensity of recycling (roughly 5% of primary production) yields electricity cost savings during peak months that are insufficient to offset the depreciation of idle capacity; and (3) the temporal rigidity of scrap supply and its high inventory holding costs render seasonal shutdowns economically unviable. By exploring various demand scenarios, we evaluate how the value of overcapacity evolves as the industry transitions from a primary-dominated era toward a recycling-dominated one.
The flexibility value of aluminum smelting and its resulting operational patterns are inherently governed by the physical constraints of the smelting process, specifically the extent to which individual potlines can adjust their power consumption and the technical difficulty associated with complete shutdowns and restarts. Using performance data from numerical simulations of aluminum smelter operations, as well as existing literature on the shutdown, restart processes, and costs of real-world aluminum smelters, we developed a reduced-order model that represents the key characteristics of aluminum electrolytic cell temperature constraints and restart costs (see Supplementary Note 3 for more details). However, the cost uncertainty is large because there are few historical cases of aluminum smelters being restarted. Thus, we evaluated four comparative scenarios:
@@
Finally, we introduce a scenario dimension concerning labor and employment flexibility, which may affect the economic feasibility of seasonal operation. We evaluate two comparative cases: (1) Inflexible labor (Core scenario): This represents the most unfavorable case where smelters must maintain their full workforce and associated labor and standby costs regardless of output levels or seasonal shutdowns. Workers are assumed to be unable to transfer to other industries. (2) Flexible labor: This case explores the potential for labor mobility and further cost optimization. If a potline is shut down, labor and standby costs are reduced. By modeling combinations of different scenarios for primary aluminum demand, smelter flexibility, technology costs, and labor flexibility, we are able to evaluate the sensitivity of the main power system and aluminum industry outcomes to changes along these dimensions.
4.2 Aluminum demand projections
Our objective is to project China’s primary aluminum demand for the coming decades, which requires projecting both total aluminum demand and the future supply of recycled aluminum. Our methodology is based on the assumption that per capita aluminum in-use stock tends to saturate as the economy develops, a common approach for projecting the demand for metal raw materials like steel and aluminum [51, 31]. To obtain historical aluminum in-use stock data, which is not explicitly tracked in official statistics, we sourced historical aluminum production data from the China Statistical Yearbook. We then calculated the aluminum stock by accumulating historical production and accounting for product retirement. Specifically, using China’s primary and secondary aluminum production data from 1993 to 2023 and assuming that the lifespan of aluminum products follows a Weibull distribution, we calculated the annual aluminum in-use stock up to 2024. Next, we input this data into the EnergyPathways (EP) model [17], a mature, open-source software for forecasting commodity demands in energy systems, which has been used in high-profile studies such as the Net-Zero America [28] and Net-Zero Australia [13] projects. The choice of EP is motivated by its proven capability to model future product demand. Using the stock-based saturation method and projected future population from Chen et al. [7], we used logistic regression (S-curve fitting) to extrapolate future aluminum demand. This approach is well-suited for our study as it captures the decline in primary metal consumption due to slowing population growth, despite continuing economic growth—a critical feature for analyzing the upcoming overcapacity in China’s aluminum sector as it matures. The EP model projects the annual aluminum stock, total aluminum demand, and retired aluminum products (the available scrap in that year) up to 2060. The calculation formula is as follows:
@@ 4.2 Aluminum demand projections
4.3 Modeling smelter operational flexibility
This study employs a mixed-integer linear model, an initial version of which was developed by Shen et al. [47] as a linear model for production and energy consumption decisions in smelter operations. The model is capable of representing the relationships between internal temperature changes and electrolysis rates in response to applied voltage and current in a potline, as observed in both numerical simulations and experiments [41, 42]. It is transformed into a linear programming model through piecewise linearization and numerical approximation. To capture the unique flexibility characteristics of aluminum smelters, we introduce binary variables for startup and shutdown decisions, which extends this framework to a mixed-integer linear programming (MILP) model. This extension is essential because aluminum smelters, similar to thermal power units, exhibit discrete operational states and path-dependent costs. Specifically, if a potline remains operational, it must maintain a minimum power level (typically 70–90% of rated power) to preserve the internal thermal balance of the electrolytic cells. Reducing power further requires a complete shutdown, which then incurs significant restart costs to restore the thermal and chemical equilibrium—costs that can be orders of magnitude higher than those for conventional power units [48]. A pure linear model would fail to capture these ”all-or-nothing” and ”minimum-stable-output” constraints, potentially overestimating the smelter’s ability to provide continuous, low-cost load shedding. This enables our model to optimize the smelter’s operational decisions over a longer timescale (see Supplementary Note 3 and 5). However, the introduction of integer variables makes it computationally infeasible to directly embed a plant-level operational model into a power system planning model. Therefore, we design a decomposition-based iterative algorithm for solving this problem. To determine the production capacity of smelters, based on the Chinese government’s policy that primary aluminum capacity will not exceed 45 million tonnes per year, we assume that the production capacity of aluminum smelters remains unchanged at the 2024 level of 45 million tonnes per year, unless early retirement is chosen. This assumption is based on the typical lifespan of an aluminum smelter, which exceeds 50 years, and the fact that more than 90% of China’s aluminum smelters were built after 2000 (see Supplementary Note 3.3).
@@ 4.4 Smelter costing
4.5 Power system capacity expansion model
@@
We utilize the Python for Power System Analysis (PyPSA) framework, a widely-recognized, open-source tool for high-resolution energy system optimization. PyPSA has been extensively applied in numerous peer-reviewed studies to explore large-scale decarbonization pathways, most notably within the European context through the PyPSA-Eur project [22]. Unlike “copper plate” models, PyPSA captures the essential physics of transmission through a linearized DC approximation of power flow, enforcing Kirchhoff’s laws and identifying network bottlenecks that drive storage requirements and renewable curtailment. Furthermore, PyPSA natively handles sector coupling, allowing electricity, heat, hydrogen, and gas to be co-optimized—a requirement for capturing the interaction between industrial electrification and electricity sector dynamics. In this study, we utilize PyPSA-China, a peer-reviewed model build on PyPSA that incorporates the specialized dataset for China’s net-zero transition analysis [65].
4.6 Dimensionality reduction and tailored iterative solution
@@
Our objective is to address the computational intractability that arises when integrating detailed smelter operation models (which include integer variables) into a large-scale power system planning model. Embedding millions of system variables alongside thousands of discrete industrial constraints would result in a “curse of dimensionality” that is numerically unsolvable. We adopt an iterative decomposition framework based on the Data-Driven Dimension Reduction (D3R) method proposed by Lyu et al. [37]. Unlike traditional analytical approximations that rely on pre-specified geometric templates (e.g., box or ellipsoid constraints), D3R utilizes inverse optimization to “learn” the feasible operating region of complex industrial loads. It trains a low-dimensional polytope that best approximates the smelter’s physical flexibility (including thermal and stoichiometry constraints) while remaining compatible with linear optimization. The D3R framework has been rigorously verified for its ability to maintain high approximation accuracy while reducing the computational burden by orders of magnitude, making it uniquely suitable for modeling the flexibility potential of energy-intensive industries in large-scale energy systems [37].
@@
The process begins by embedding this linearized smelter model obtained through the D3R framework into the main linear capacity expansion model. Solving this initial integrated model yields a first estimate of electricity system capacity expansion results and the hourly local marginal prices (LMPs) for electricity in each province. Then, these LMPs are fixed and used as an exogenous price signal for the detailed smelter operational models. This step, a core part of the decomposition, allows each smelter’s 8760-hourly cost-minimization problem (including all unit commitment constraints) to be solved independently and in parallel. Once the detailed optimal operational schedule (i.e., the 8760-hourly load profile) is determined for each smelter, we fix these schedules in the main capacity expansion model and then re-optimize all remaining power system planning and operational variables. This re-optimization generates a new, updated set of LMPs, which are fed back into the detailed smelter models. This entire process is iterated until the objective function of the planning model converges—which our tests show occurs within several iterations—indicating a consistent, optimal solution for both system planning and smelter operation has been found. For more details, see Supplementary Note 5.
4.7 Limitations and opportunities

### 17 Apr 2026 v3

# Industrial overcapacity can enable seasonal flexibility in electricity use

@@ Abstract
In many countries, declining demand in energy-intensive industries (EIIs) such as cement, steel, and aluminum is leading to industrial overcapacity. Although industrial overcapacity is traditionally envisioned as problematic and resource-wasteful, it could unlock EIIs’ flexibility in electricity use. Here, using China’s aluminum smelting industry as a case study, we evaluate the system-level cost-benefit of retaining EII overcapacity for flexible electricity use in decarbonized energy systems. We find that overcapacity can enable aluminum smelters to adopt a seasonal operation paradigm, ceasing production during winter load peaks that are exacerbated by heating electrification and renewable seasonality. This seasonal operation paradigm could reduce the investment and operational costs of China’s decarbonized electricity system by 23-32 billion CNY/year (11-15% of the aluminum smelting industry’s product value), sufficient to offset the increased smelter maintenance and product storage costs associated with overcapacity. It may also create labor complementarities between the aluminum and thermal power sectors.
@@ Introduction
In this paper, we aim to understand the value of aluminum smelting overcapacity for providing demand flexibility in China’s decarbonized energy system. We examine these effects across both physical infrastructure and socio-economic dimensions (e.g., labor deployment), as both are essential for charting a sustainable path for the aluminum industry. However, the computational complexity of modeling and optimizing the intricate coupling between two already large and complex systems—the electricity system and aluminum smelters— is a significant barrier to such analysis and has not been done before. We developed an integrated provincial-level power-smelting co-optimization model (Figure 1). This approach couples the PyPSA power system framework [6] with a detailed aluminum smelter operation model [48] for hourly co-optimization. To solve this computationally heavy 8760-hour simulation, we adopted a Data-Driven Dimension Reduction (D3R) framework [37] paired with an iterative decomposition algorithm. First, D3R linearizes complex smelter constraints and generates an initial co-optimized solution to make the power system model easier to solve. The power system model then passes hourly locational marginal prices (LMPs) to the smelters. Using these price signals, each smelter independently optimizes its discrete production decisions (e.g., shutdown and restart costs) in parallel. This price-load feedback loop iterates until convergence, bridging modeling high-fidelity smelting physics with large-scale grid planning (see Methods for details).
@@
Overcapacity in aluminum smelting in the recycle era
Figure 1: Schematic diagram of the integrated power-aluminum co-optimization framework. To overcome the computational intractability of an 8760-hour horizon, the PyPSA-based grid model and detailed smelter operations are decoupled and solved iteratively through the exchange of local marginal prices (LMPs) and industrial load profiles, aided by the Data-Driven Dimension Reduction (D3R) framework.
@@
The restart-up costs are measured from the electricity input side. Exchange rates throughout the study (annual averages 2020): 1 USD = 6.900 CNY; 1 EUR = 7.868 CNY; 1 USD = 0.877 EUR. The electricity system cost data is primarily from the PyPSA-China model [64], with the specific Chinese cost assumptions and their original sources documented in Supplementary Table 7. Long-duration storage includes H2 and pumped hydro storage. Our core scenario is defined by the combination of Mid primary aluminum demand, Mid smelter operational flexibility, Mid technology costs, and Inflexible labor.
This study employs a scenario analysis approach to explore the impact of key uncertainties on the system-level value of aluminum smelting overcapacity. As detailed in Table 1, we evaluate four primary dimensions: (i) primary aluminum demand, which determines the industry’s aggregate flexibility potential by defining the size of the sector; (ii) smelter operational flexibility, representing the physical minimum load constraints and the economic penalty (e.g., restart costs) of non-baseload operation; (iii) technology costs, reflecting the relative competitiveness of demand-side flexibility against supply-side flexibility alternatives (e.g., batteries and hydrogen); and (iv) labor flexibility, which dictates whether workforce costs can be mitigated during seasonal shutdowns. (see Methods for more details) To ensure the robustness and traceability of these scenario settings, the bounds for each dimension are systematically derived from empirical literature and industry standards. For technology costs, the Mid scenario adopts peer-reviewed trajectories from the PyPSA-China[64] dataset, while the Low and High bounds for the costs are scaled by 20-20% and +50+50%, respectively, adhering to the AACE International Class 4 estimate guidelines [4].
@@
Electricity system savings through flexible smelting
@@
Cost-benefit analysis of retaining smelting overcapacity
@@
Figure 4: Trade-off analysis and cost-effective strategy for retaining overcapacity. In (a), electricity system cost savings, smelter cost increase, and net system benefit with different retained capacities in 2050 in the core scenario are shown. The maximal net benefit (electricity system cost reduction minus smelter increased costs) is achieved with 30% overcapacity in 2050. In (b), the mean optimal overcapacity rate increases from 6% in 2030 to 30% in 2050, due to higher flexibility requirements as the electricity system approaches net-zero emissions. Notably, while the optimal rate of overcapacity rises, the absolute cost-effective smelting capacity decreases over time due to declining primary aluminum demand. The net benefit across different scenarios first decreases as primary aluminum demand declines to 2040 levels, then increases in the net-zero emissions scenario in 2050, where load flexibility is more valuable. All statistics are derived from $n=36$ independent scenarios per planning year. For each box plot, the center horizontal line represents the median, while the top and bottom box limits indicate the third and first quartiles, respectively.
@@
Seasonal operation paradigm enabled by overcapacity
@@
(b) Monthly electricity demand and generation in 2050. (Residual demand = electricity demand - electricity generation)
Figure 5: Seasonal operation of aluminum smelters is complementary to energy system seasonality. In (a), smelter production largely ceases from mid-November to mid-March with demand met by stored inventory, but operates at full capacity from late March to early November. Comparing (a) with (b), the electricity consumption of aluminum smelting is largely complementary to China’s monthly residual load, acting as a crucial seasonal flexibility resource to mitigate the winter electricity demand peaks driven by heating electrification.
@@
Panel (b) reveals the underlying system driver for this behavior: the emergence of a “seasonal duck curve”. Driven by high renewable penetration and heating electrification, China’s residual demand (electricity demand minus electricity generation) exhibits a distinct and prolonged winter peak. During these winter periods, renewable energy output, especially from hydropower, is at its minimum while heating demand peaks. Although wind power also increases in the winter, it is outpaced by the surge in heating demand, forcing the power system to rely more heavily on gas and coal generation to ensure power supply reliability. Here, the national aggregate solar output obscures regional differences (e.g., reduced summer solar generation in Yunnan due to the monsoon season). For provincial results, see Supplementary Figure 8.
Figure 6: Comparison of the change in levelized cost per tonne of aluminum between the $20$ baseline and various 2050 overcapacity levels. In the core scenario, maintaining 30% overcapacity is beneficial for the smelters, reducing aluminum production costs by over 1,500 CNY/tonne (approximately 9% of current production costs), with the drop in electricity bills offsetting the maintenance costs of overcapacity and increased restart, storage, and capital costs.
@@ Discussion
Realizing this seasonal paradigm also requires addressing its socio-economic implications, particularly concerning labor. However, because the labor intensity of modern primary aluminum production is exceptionally low (accounting for only 2%–3% of total costs [46]), the immense energy cost savings achieved during high-generation periods are more than sufficient to cover standby wages or training stipends during winter lulls. As financial incentives make the transition to seasonal operations economically compelling, exploring potential employment complementarity under this paradigm represents a promising avenue for future research and policy design. For instance, the winter production lull in aluminum smelting naturally coincides with the operational peaks of thermal power plants in a decarbonized grid. While facilitating such inter-industry job mobility requires specific spatial planning (e.g., large integrated industrial parks) and targeted institutional support, it is well-grounded in China’s empirical reality. Formalized national policies such as the “shared employee” framework, alongside the internal labor markets of large state-owned conglomerates, already facilitate similar seasonal workforce rotations across manufacturing and heavy industries (see Supplementary Notes 6.1). Ultimately, supported by carefully designed market and policy mechanisms, the seasonal operation of energy-intensive industries holds the potential to unlock broader societal benefits, transforming structural overcapacity into a strategic asset for employment resilience and social stability.
Methods
Data Source and Scenario Settings
@@ Methods
In our model, China’s domestic aluminum smelters are set to meet the annual primary aluminum demand, defined as the total demand minus the availability of recycled aluminum. This primary demand is a critical factor determining the system-level value of flexibility and the optimal level of strategically retained overcapacity: intuitively, a higher primary demand implies a larger industrial base with greater cumulative potential for seasonal load shifting and peak shaving. We focus on primary aluminum smelters for seasonal flexibility, modeling recycled (secondary) aluminum production as a continuous baseload process. This constraint is driven by three techno-economic factors: (1) unlike the legacy overcapacity in primary smelting, the rapidly expanding recycling sector requires substantial new capital investment, necessitating high utilization rates to amortize marginal capital expenditures; (2) the low energy intensity of recycling (roughly 5% of primary production) yields electricity cost savings during peak months that are insufficient to offset the depreciation of idle capacity; and (3) the temporal rigidity of scrap supply and its high inventory holding costs render seasonal shutdowns economically unviable. By exploring various demand scenarios, we evaluate how the value of overcapacity evolves as the industry transitions from a primary-dominated era toward a recycling-dominated one.
The flexibility value of aluminum smelting and its resulting operational patterns are governed by the physical constraints of the smelting process, specifically the extent to which individual potlines can adjust their power consumption and the technical difficulty associated with complete shutdowns and restarts. Using performance data from numerical simulations of aluminum smelter operations, as well as existing literature on the shutdown, restart processes, and costs of real-world aluminum smelters, we developed a reduced-order model that represents the key characteristics of aluminum electrolytic cell temperature constraints and restart costs (see Supplementary Note 3 for more details). However, the cost uncertainty is large because there are few historical cases of aluminum smelters being restarted. Thus, we evaluated four comparative scenarios:
@@
Finally, we introduce a scenario dimension concerning labor and employment flexibility, which may affect the economic feasibility of seasonal operation. We evaluate two comparative cases: (1) Inflexible labor (core scenario): This represents the most unfavorable case where smelters must maintain their full workforce and associated labor and standby costs regardless of output levels or seasonal shutdowns. Workers are assumed to be unable to transfer to other industries. (2) Flexible labor: If a potline is shut down, labor and standby costs are reduced. By modeling combinations of different scenarios for primary aluminum demand, smelter flexibility, technology costs, and labor flexibility, we are able to evaluate the sensitivity of the main power system and aluminum industry outcomes to changes along these dimensions.
Aluminum demand projections
Our objective is to project China’s primary aluminum demand for the coming decades, which requires projecting both total aluminum demand and the future supply of recycled aluminum. Our methodology is based on the assumption that per capita aluminum in-use stock tends to saturate as the economy develops, a common approach for projecting the demand for metal raw materials like steel and aluminum [50, 30]. To obtain historical aluminum in-use stock data, we sourced historical aluminum production data from the China Statistical Yearbook. We then calculated the aluminum stock by accumulating historical production and accounting for product retirement. Specifically, using China’s primary and secondary aluminum production data from 1993 to 2023 and assuming that the lifespan of aluminum products follows a Weibull distribution, we calculated the annual aluminum in-use stock up to 2024. Next, we input this data into the EnergyPathways (EP) model [17], a mature, open-source software for forecasting commodity demands in energy systems, which has been used in high-profile studies such as the Net-Zero America [28] and Net-Zero Australia [13] projects. The choice of EP is motivated by its proven capability to model future product demand. Using the stock-based saturation method and projected future population from Chen and others [7], we used logistic regression (S-curve fitting) to extrapolate future aluminum demand. This approach is well-suited for our study as it captures the decline in primary metal consumption due to slowing population growth, despite continuing economic growth. The EP model projects the annual aluminum stock, total aluminum demand, and retired aluminum products (the available scrap in that year) up to 2060. The calculation formula is as follows:
@@
Modeling smelter operational flexibility
This study employs a mixed-integer linear model, an initial version of which was developed by Shen et al. [48] as a linear model for production and energy consumption decisions in smelter operations. The model is capable of representing the relationships between internal temperature changes and electrolysis rates in response to applied voltage and current in a potline, as observed in both numerical simulations and experiments [41, 42]. It is transformed into a linear programming model through piecewise linearization and numerical approximation. To capture the unique flexibility characteristics of aluminum smelters, we introduce binary variables for startup and shutdown decisions, which extends this framework to a mixed-integer linear programming (MILP) model. This extension is essential because aluminum smelters, similar to thermal power units, exhibit discrete operational states and path-dependent costs. Specifically, if a potline remains operational, it must maintain a minimum power level (typically 70–90% of rated power) to preserve the internal thermal balance of the electrolytic cells. Reducing power further requires a complete shutdown, which then incurs significant restart costs to restore the thermal and chemical equilibrium—costs that can be orders of magnitude higher than those for conventional power units [49]. A pure linear model would fail to capture these all-or-nothing and minimum-stable-output constraints, potentially overestimating the smelter’s ability to provide continuous, low-cost load shedding. This enables our model to optimize the smelter’s operational decisions over a longer timescale (see Supplementary Note 3 and 5). However, the introduction of integer variables makes it computationally infeasible to directly embed a plant-level operational model into a power system planning model. Therefore, we design a decomposition-based iterative algorithm for solving this problem. To determine the production capacity of smelters, based on the Chinese government’s policy that primary aluminum capacity will not exceed 45 million tonnes per year, we assume that the production capacity of aluminum smelters remains unchanged at the 2024 level of 45 million tonnes per year, unless early retirement is chosen. This assumption is based on the typical lifespan of an aluminum smelter, which exceeds 50 years, and the fact that more than 90% of China’s aluminum smelters were built after 2000 (see Supplementary Note 3.3).
@@
Power system capacity expansion model
@@
We utilize the Python for Power System Analysis (PyPSA) framework, a widely-recognized, open-source tool for high-resolution energy system optimization. PyPSA has been extensively applied in numerous peer-reviewed studies to explore large-scale decarbonization pathways, most notably within the European context through the PyPSA-Eur project [22]. Unlike copper-plate models, PyPSA captures the essential physics of transmission through a linearized DC approximation of power flow, enforcing Kirchhoff’s laws and identifying network bottlenecks that drive storage requirements and renewable curtailment. Furthermore, PyPSA natively handles sector coupling, allowing electricity, heat, hydrogen, and gas to be co-optimized. In this study, we utilize PyPSA-China, a peer-reviewed model build on PyPSA that incorporates the specialized dataset for China’s net-zero transition analysis [64].
Dimensionality reduction and tailored iterative solution
@@
Our objective is to address the computational intractability that arises when integrating detailed smelter operation models (which include integer variables) into a large-scale power system planning model. We adopt an iterative decomposition framework based on the Data-Driven Dimension Reduction (D3R) method proposed by Lyu et al. [37]. D3R utilizes inverse optimization to learn the feasible operating region of complex industrial loads. It trains a low-dimensional polytope that best approximates the smelter’s physical flexibility (including thermal and stoichiometry constraints) while remaining compatible with linear optimization. The D3R framework has been rigorously verified for its ability to maintain high approximation accuracy while reducing the computational burden by orders of magnitude, making it uniquely suitable for modeling the flexibility potential of energy-intensive industries in large-scale energy systems [37].
@@
The solution process begins by embedding this linearized smelter model obtained through the D3R framework into the electricity system capacity expansion model. Solving this initial integrated model yields a first estimate of electricity system capacity expansion results and the hourly local marginal prices (LMPs) for electricity in each province. Then, these LMPs are fixed and used as an exogenous price signal for the detailed potline-level smelter operation models. This step allows each smelter’s 8760-hourly cost-minimization problem (including all unit commitment constraints) to be solved independently and in parallel. Once the detailed optimal production schedule and therefore the 8760-hourly load profile is determined for each smelter, we fix these schedules in the main capacity expansion model and then re-optimize all remaining power system expansion and dispatch variables. This re-optimization generates a new, updated set of LMPs, which are fed back into the detailed smelter models. This entire process is iterated until the objective function of the planning model converges. For more details, see Supplementary Note 5.
Limitations and opportunities